solve by factoring
x^2 – 13x = -12

Answer by nerdybill(7384)


x^2 – 13x = -12
.
begin by adding 12 to both sides to get:
x^2 – 13x + 12 = 0
.
Next, you basically are looking for "factors" of 12 that when combined will give you the -13. In this case, it would be -1 and -12:
.
(x-1)(x-12) = 0
.
Now, to find the solution, set each of the factor to zero and solve:
x-1 = 0
x = 1
and
x-12 = 0
x = 12
.
Solution: x = {1,12}